OpenStudy (anonymous):

A line has a rise of 6 and a slope of 1/20. What is the run? 30 90 120

OpenStudy (unklerhaukus):

\[\text{slope} = \frac{\text {rise} }{\text {run}}\] \[\rightarrow\text{run}=\frac{\text {rise}}{\text{slope}}\]\[=\frac{6}{1/20}\]

OpenStudy (anonymous):

so 120

OpenStudy (unklerhaukus):

checking \[\frac{6}{120}=\frac{1}{20}=\text{slope}\] you have found the correct solution